Find the missing coordinate to complete the ordered pair so that the point
( , 7) is on the line 4x + y = - 17

Answer by jim_thompson5910(35256)

4x + y = - 17

4x + 7 = - 17

4x = - 17-7

4x = - 24

x = - 24/4

x = -6


So the missing coordinate is -6 and the ordered pair that lies on the line is (-6, 7)